OpenStudy (samigupta8):

The differential representing the family of curves y^2=2c(x+√c) where c>0 is a parameter is of order and degree??? My ans is order-2 n degree-1

OpenStudy (samigupta8):

Bt it's wrong ....How?

OpenStudy (phi):

what is the differential equation?

OpenStudy (samigupta8):

If u differentiate it u will get a differential equation....

Parth (parthkohli):

Does this mean something like\[y^2=2c(x+\sqrt c)\]is a solution to some differential equation, and we have to find that?

OpenStudy (faiqraees):

order I recall is the order of the derivative of the highest order present in the equation is the order of the equation

OpenStudy (samigupta8):

Yep @parthkohli it's the solution which is given and we need the differential equation

OpenStudy (phi):

yes, but what is the specific differential equation for the given curve

Parth (parthkohli):

Oh, I can kind of see the differential equation in this because of terms like \(y^2/2\) and \(c^{3/2}\). Again, I've not studied this but let me try to come up with something, yeah.

OpenStudy (samigupta8):

Well I always differentiate the equation given untill i eliminate the arvitary constant in it...

OpenStudy (samigupta8):

Sorry arbitary*

OpenStudy (samigupta8):

Like if u differentiate the expression first time u will get 2ydy/dx=2c Second time diff will yield Yd2y/dx^2 +(dy/dx)^2=0

OpenStudy (phi):

2ydy/dx=2c I think they are asking about the order and degree of this equation

OpenStudy (samigupta8):

Nope , we nees to remove the arbitrary constants in order to get the differential equation

Parth (parthkohli):

Well, I got this:\[y'' = -\frac{y'^2}{y}\]

OpenStudy (samigupta8):

Need*

OpenStudy (samigupta8):

Yep parth...this is correct that's what i wrote above...

Parth (parthkohli):

Oh great, exactly the same.

OpenStudy (samigupta8):

From this v can see that highest order of differential is 2 n it's degree is 1

Parth (parthkohli):

Yeah, I guess you're right.

OpenStudy (samigupta8):

Bt the ans is very different from what v r getting. .

Parth (parthkohli):

Does the degree of a DE reflect the highest power of the variable, or can it also include the powers of the derivatives?

OpenStudy (faiqraees):

what is the answer?

OpenStudy (samigupta8):

Order -1 , degree -3

OpenStudy (samigupta8):

Okk well @parthkohli order is the highest order derivative ...n degree is the order's power....

OpenStudy (samigupta8):

Likewise in the present situation we r having y" as the highest derivative n so order wud be 2 n it's power is 1 so that turns out to be the degree of differential

OpenStudy (rational):

One arbitrary constant in the general solution implies that the order of the corresponding differential equation is exactly 1.

OpenStudy (samigupta8):

Yep @rational bt here we r having two arbitary constants....

OpenStudy (rational):

I see only one arbitrary constant, c ?

OpenStudy (samigupta8):

If you open the brackets u will get Y^2=2cx + c√c Which can be rewritten as Y^2=2cx + k Here c n k are two arbitrary constant

OpenStudy (samigupta8):

Okk sorry 2c√c....

OpenStudy (rational):

\(k\) is not arbitrary if it depends on \(c\)

OpenStudy (samigupta8):

Okk....bt what about the differential we formed over here ....as per that we r getting order as 2...

Parth (parthkohli):

Hey rational, I wonder... is there a nice way to define "order" of a differential equation? Like do we just call it the most efficient way to express a differential equation?

OpenStudy (samigupta8):

Bt c is also arbitary then if any parameter is arbitary n some other thing depends on it then it is also arbitrary isn't it?

OpenStudy (rational):

Yes, but you don't have two different arbitrary constants. You get to control only one value which changes the value at both places

OpenStudy (rational):

Hey @ParthKohli compare this with the difference equations

OpenStudy (rational):

or recurrence relations

OpenStudy (rational):

How many initial values do you need to know to solve the particular solution of the following recurrence relation ? \[a_{n+2} = a_{n+1}+a_n\] ?

OpenStudy (samigupta8):

Yep exactly @rational i agree with you... If we put the value of c to be something then that wud imply the change in value of k as well...

Parth (parthkohli):

Two of them. I agree. But what if we're specified with a differential equation with a huge order, and we want to know if it's the best way of being expressed?

Parth (parthkohli):

In a way, I cannot claim that \(x = 1\) is a better way of expressing \(x^2 = 1\) so my question is kinda stupid. Is it the same with DE?

OpenStudy (rational):

Yes, The differential equations, \(y' = 2\) and \(y''=0\) model different problems.

Parth (parthkohli):

Ah, nice. Do these also model different problems?\[y'' = -y\]\[y''' = -y'\]

OpenStudy (rational):

Ofcourse, they are different

OpenStudy (rational):

First equation models block spring oscillator without damping

OpenStudy (rational):

if i remember correctly..

OpenStudy (samigupta8):

Yep...

OpenStudy (rational):

second equation may model some other situation with some damping... i don't know...

Parth (parthkohli):

Hmm, but seems that the only difference between these two DEs is an added arbitrary constant. So they're not as different as, say, \(y' = 2\) and \(y'' = 0\). Well, alright, whatever I'm saying is subjective. lol

OpenStudy (samigupta8):

Bt @rational can we differentiate the equation infinite times to get the order of the differential.. In such a way we can have any number of possibility of order ..

OpenStudy (samigupta8):

I were taught that we have to jst eliminate the arbitary constants that's it !!!!!!

OpenStudy (rational):

\(y''' = -y'\) can be thought of as the following system : \(y' = x\) \(x''=-x\)

OpenStudy (rational):

Let me think of a better reason to see them as different...

Parth (parthkohli):

Nah, I totally understand now. Thanks :)

OpenStudy (rational):

@samigupta8 in general, yes. In our specific problem, we're given the number of arbitrary constants. So order is fixed...

OpenStudy (samigupta8):

Bt it is always like that we would be given with a specific problem and which would contains some arbitrary constant and the order is going to be their count....

OpenStudy (rational):

Let me ask you a question

OpenStudy (samigupta8):

N @rational shall we not follow the procedure as mentioned by @parthkohli and me which involves differentiation of the given curve to get differential

OpenStudy (samigupta8):

Yep sure....

Parth (parthkohli):

\[y y' = c \]Now\[y^2 = yy' x + y^{3/2} y'^{3/2}\]\[ (y^2 -yy'x)^2 = y^3 y'^3 \]Well, I'm not sure again, because all this degree stuff is pretty sensitive.

OpenStudy (rational):

Solve the following initial value problem with \(y(0) = \sqrt{2}\): \[y' = 1/y\]

OpenStudy (samigupta8):

Y^2=2x+2

OpenStudy (rational):

Excellent!

Parth (parthkohli):

Oh, and if you hate that \(x\) in the equation substitute something like\[\frac{y^2}{2c} -\sqrt c=x\]

OpenStudy (samigupta8):

Bt what does this prove ???

OpenStudy (samigupta8):

Well now i m getting confused over the arbitrary constant even? @rational would u please clarify it....That how to check for the count of arbitrary constant in a particular equation

OpenStudy (anonymous):

\[y^2=2c \left( x+\sqrt{c} \right)\] diff.w.r.t. x \[2y \frac{ dy }{ dx }=2c*1,c=y \frac{ dy }{ dx },\sqrt{c}=\sqrt{y \frac{ dy }{ dx }}\] diff. eq. is \[y^2=2y \frac{ dy }{ dx }\left( x+\sqrt{y \frac{ dy }{ dx }} \right)\] \[y=2\frac{ dy }{ dx }\left( x+\sqrt{y \frac{ dy }{ dx }} \right)\]

OpenStudy (samigupta8):

After finding the value of c to be dy/dx why didn't u differentiate it again in order to get differential..?

OpenStudy (priyar):

surjithayer is right.. y=2y'(x+(yy')^1/2 ) is the differential eq... here u can see that the order is 1. and on squaring on both sides.. u can see that the degree will be three..

OpenStudy (priyar):

also u can say that the order is 1 just by looking at the diff eq. as the order is equal to the no.of times we must differentiate and that is equal to the no. of. arbitary constants...here there is just "1" arbitary constant which is "c"..

OpenStudy (samigupta8):

Bt why didn't u differentiate it again to get the differential???

OpenStudy (priyar):

u should only do the minimum diff. required and thats equal to the no. of. arbitrary constants..

OpenStudy (anonymous):

here only one arbitrary constant so differentiate only once.